Given the demand and supply function for the three goods





qd1=20-p1-p3 ;qs1=-10+p1





qd2=40-2p2-p3 ;qs2=2p2





qd3=10-p1+p2-p3 ;qs3=-5+3p3



A Write the equilibrium condition?



B find the equilibrium price?




1
Expert's answer
2022-01-17T09:55:44-0500

"a)"

Equilibrium in the three commodity market occur when the supply for each of the 3 commodities equal to its demand.

That is,

"Qd_1=Qs_1\\Qd_2=Qs_2\\Qd_3=Qs_3"

This is the equilibrium condition.


"b)"

To find the equilibrium price we proceed as follows.

For commodity 1,

"Qd_1=Qs_1"

So,

"20-p_1-p_3=-10+p_1\\

2p_1+p_3=30......(1)"


For commodity 2,

"Qd_2=Qs_2"

So,

"40-2p_2-p_3=2p_2\\

4p_2+p_3=40...(2)"


For commodity 3,

"Qd_3=Qs_3"

So,

"10-p_1+p_2-p_3=-5+3p_3\\

p_1-p_2+4p_3=15.....(3)"


Equation (1)-Equation (2) gives,

"2p_1-4p_2=-10.....(4)"


"Equation(3)-(4\times Equation(2))" gives,

"p_1-17p_2=-145.....(5)"


Solving equation (4) and (5),

"Equation(4)-(2\times Equation(5))\\

30p2=280\implies

p2={28\over3}"


From equation (4), we can obtain the value of "p_1"

"2p_1-(4\times{28\over3})=-10\\

2p_1={82\over3}\implies

p1={41\over3}"


From equation (1) we can obtain the value "p_3"

So,

"(2\times{41\over3})+p_3=30\\

p_3=30-{82\over3}\\
Therefore,\\

p_3={8\over3}"


Therefore, the equilibrium prices for commodities 1,2,3 are "p_1={41\over3}, p_2={28\over3}, and\space p_3={8\over3}" respectively.